(1 point) A. Determine the sample size required to estimate a population proportion to within 0.032 with 99% confidence, assuming that you have no knowledge of the approximate value of the sample proportion.

Sample Size =  

B. Repeat the previous problem, but now with the knowledge that a prior study found a sample proportion of approximately 0.33.

Sample Size =  

Note: The table in Sullivan that gives values for ??/2zα/2 is not accurate. Make sure to use the following values for ?z.

?0.10z0.10 1.282
?0.05z0.05 1.645
?0.025z0.025 1.960
?0.01z0.01 2.326
?0.005z0.005 2.576
Answer #1

Solution :

Given that,

= 0.33

1 - = 0.67

margin of error = E = 0.032

Z/2 = 2.576

sample size = n = (Z / 2 / E)2 * * (1 - )

= (2.576 / 0.032)2 * 0.33 * 0.67

= 1433

sample size = n = 1433